Chlorine can be prepared in the laboratory by the reaction of manganese dioxide with hydrochloric acid, HCl(aq), as described by the chemical equation


MnO2(s)+4HCl(aq)⟶MnCl2(aq)+2H2O(l)+Cl2(g)

How much MnO2(s) should be added to excess HCl(aq) to obtain 245 mL Cl2(g) at 25 °C and 715 Torr?


1
Expert's answer
2021-11-12T10:01:01-0500

Molar mass of Cl2= 70.9

= 245/70.9= 3.46moles

Molar mass of MnO2= 86.94

= 3.46×86.94=300.43g